### v1

# Key Words.: Gravitational lensing: strong – Methods: data analysis – Methods: observational – Galaxies: clusters: individual: Perseus1 Introduction

Euclid: The Early Release Observations Lens Search Experiment††thanks: This paper is published on behalf of the Euclid Consortium.
We investigate the ability of the Euclid telescope to detect galaxy-scale gravitational lenses. To do so, we perform a systematic visual inspection of the 0.7 deg2 Euclid ERO data towards the Perseus cluster using both the high-resolution VIS band, and the lower resolution NISP bands. We inspect every extended source brighter than magnitude 23 in with 41 expert human classifiers. This amounts to 12 086 stamps of . We find 3 grade A and 13 grade B candidates. We assess the validity of these 16 candidates by modelling them and checking that they are consistent with a single source lensed by a plausible mass distribution. Five of the candidates pass this check, five others are rejected by the modelling and six are inconclusive. Extrapolating from the five successfully modelled candidates, we infer that the full of the Euclid Wide Survey should contain galaxy-galaxy lenses that are both discoverable through visual inspection and have valid lens models. This is consistent with theoretical forecasts of discoverable galaxy-galaxy lenses in Euclid. Our five modelled lenses have Einstein radii in the range , but their Einstein radius distribution is on the higher side when compared to theoretical forecasts. This suggests that our methodology is likely missing small Einstein radius systems. Whilst it is implausible to visually inspect the full Euclid data set, our results corroborate the promise that Euclid will ultimately deliver a sample of around galaxy-scale lenses.
Key Words.:
Gravitational lensing: strong – Methods: data analysis – Methods: observational – Galaxies: clusters: individual: Perseus1 Introduction
Strong gravitational lensing by massive galaxies offers a plethora of applications both in cosmology and astrophysics. Some notable examples include measuring the total mass of lens galaxies within the Einstein radius, as well as disentangling the contributions from visible and dark matter components (e.g., Auger et al. 2009). When coupled with deep spectroscopic observations, it allows for the placement of constraints on the stellar initial mass function of lens galaxies (e.g., Ferreras et al. 2010; Dutton & Treu 2014; Sonnenfeld et al. 2019). Thanks to the lensing magnification, it serves as a natural telescope to study lensed sources otherwise too faint or angularly too small to be detected (e.g., Hezaveh et al. 2013), and even to map their velocity field with unprecedented spatial resolution (e.g., Paraficz et al. 2018). By studying small-scale distortions of lensed images or arcs, it allows us to infer the presence of low-mass dark halos either in the lens or along the line of sight up to the source redshift, and in turn to study the properties of dark matter (e.g., Vegetti et al. 2010; O’Riordan et al. 2023; Gilman et al. 2024). Moreover, when the lensed source is time-variable, such as a quasar or a supernova, strong lensing offers an independent way of measuring the expansion rate of the Universe, (Refsdal 1964; Wong et al. 2020; Shajib et al. 2023; Grillo et al. 2024; Pascale et al. 2024).
Lastly, the measurement of weak lensing shear from strong lensing images has been proposed as a potential cosmological probe (Birrer et al. 2017, 2018). A minimal model for the shear that is non-degenerate with lens model parameters was derived by Fleury et al. (2021), and this quantity was shown to be measurable in mock imaging data by Hogg et al. (2023). For high-precision cosmological constraints, however, a Euclid-size data set of strong lenses will be required.
Galaxy-scale strong lensing events are rare, with about one object out of thousands showing lensing features (Oguri & Marshall 2010; Collett 2015). In addition, these systems are small in angular terms, spanning between tenths of an arcsecond to a few arcseconds on the sky, and the lensed images of the sources (arcs, rings, and multiple point sources) are often hidden in the glare of the foreground lensing galaxy. Because lenses are rare, compact, and low-contrast objects, they are best discovered in deep, sharp, wide-field surveys. Euclid is exactly that (Euclid Collaboration: Mellier et al. 2024), and is the focus of the present work. In fact, lensed QSO candidates have already been proposed from Euclid observations (Cuillandre et al. 2024b).
However, even with high-quality Euclid data, finding lenses remains challenging, not only because of the intrinsic rarity of strong lensing, but also because other non-lensing objects mimic the morphology of lenses (e.g., ring galaxies, galaxy mergers, spiral galaxies, or even random alignment). So far, the best methods to discover lenses involve various flavours of convolutional neural networks (CNNs; among many others see Petrillo et al. 2019; Jacobs et al. 2019; Stein et al. 2022; Savary et al. 2022; Rojas et al. 2022) and more recently, transformer networks (Thuruthipilly et al. 2022; Grespan et al. 2024). These networks require training sets that match the instrumental and astrophysical properties of the data very closely. Since we do not possess a number of real strong lens observations sufficient for a training set, these have to be simulated, and the effectiveness of simulations in emulating nature is limited.
Machine learning methods are effective in finding lenses, but they are limited in terms of purity and false-positive rates. When false-positive rates are below 1% (e.g., Canameras et al. 2023), they are useful to pre-select candidates among parent samples of tens or even hundreds of millions of galaxies. However, as the true – and poorly known – prevalence of lenses on the plane of the sky is very low, even a 1% false-positive rate can result in samples dominated by contaminants and unconvincing candidates. This is known as the base rate fallacy. As a result, human visual inspection is almost always necessary to fine-grain the network selection, although humans themselves sometimes have difficulty deciding on the validity of a lens candidate. This is especially true for small Einstein radius lenses.
Human visual inspection is useful on its own as a lens finding method, but is limited by the volume of data that can be inspected by a team of experts in a reasonable amount of time. However, it has the advantage that it can pick up unusual lensing configurations (e.g., Keeton et al. 2000; Orban de Xivry & Marshall 2009; Collett & Bacon 2016) that, by definition, do not appear in large numbers in simulated training sets. It is also the only way to evaluate the prevalence of lenses on the plane of the sky and therefore the expected surface density of lensing systems for a given instrument, depth, spatial resolution, and wavelength. This has been attempted a few times in the past, for example in Hubble Space Telescope (HST) imaging (e.g., Faure et al. 2008; Pawase et al. 2014; Garvin et al. 2022). Rojas et al. (2023) also evaluate the performance of visual inspection with simulated data.
In the Early Release Observations (ERO) Lens Search Experiment (ELSE), we carry out a blind visual search of galaxy-scale strong lensing systems, using some of the first data from the Euclid ESA satellite (Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2024). We focus on the Euclid ERO imaging of the Perseus galaxy cluster (Cuillandre et al. 2024a) and the only selection criterion for the extended sources to be inspected is a magnitude cut. Our search is therefore one of the broadest visual searches carried out so far in terms of pre-selection. The goals are: (1) to evaluate the performance of the Euclid telescope at finding galaxy-scale lenses; (2) to study the prevalence of strong lenses found by humans in Euclid; (3) to test the efficiency of human experts at finding lenses; and, (4) to optimise future visual inspections that will be performed on lens candidates found by the automatic pipeline in Euclid.
As the humans involved in this exercise are all lensing experts but limited in number (41 exactly), our work contrasts with more intensive citizen science searches involving much larger numbers of humans but spanning a much scarcer range in terms of expertise available (Marshall et al. 2015; More et al. 2016).
This article is organised as follows. We introduce the Euclid ERO observations of the Perseus cluster in Sect. 2. Then, we explain the methodology of the visual inspections in Sect. 3, along with the visualisation tools used for them. Afterwards, we report the results from the visual inspection in Sect. 4. Next, we present the sample of lens candidates and some initial modelling in Sect. 5. Then, in Sect. 6 we compare the sample against the literature and estimate the number of lenses that will be discovered in the Euclid Wide Survey (EWS). Finally, we conclude in Sect. 7.
2 Data
The Euclid ERO programme (Euclid Early Release Observations 2024) targeted the Perseus galaxy cluster, obtaining very deep data of the central region of the cluster in , in the broad optical filter from the visible imager (VIS) instrument (Euclid Collaboration: Cropper et al. 2024), and the three broad filters , , and in the near-infrared from the Near-Infrared Spectrometer and Photometer (NISP) instrument (Euclid Collaboration: Jahnke et al. 2024).
These data were collected during the Euclid performance verification phase in September 2023 (Cuillandre et al. 2024a). All Euclid science observations adhere to a reference observing sequences (ROS, Euclid Collaboration: Scaramella et al. 2022), which consists of four dithered exposures lasting seconds each in the filter, and four dithered exposures of seconds each in the , , and filters. Four ROSs were obtained for this field, with a total integration time of seconds in the filter and seconds in the , , and filters, achieving a depth magnitudes deeper than that of the EWS, which relies on a single ROS (Euclid Collaboration: Scaramella et al. 2022). Therefore, these exceptional data reach a point-source depth of () at , with a () full width at half maximum (FWHM), and a surface brightness limit of . We refer the reader to Cuillandre et al. (2024a) for more details on the data reduction.
The Perseus cluster data set represents the ideal data to perform a first search of gravitational lenses, and prepare for the future systematic investigations with the EWS data.
The astrometrically and photometrically calibrated imaging stacks across all four Euclid bands are accompanied by catalogues for compact sources produced using the tool SourceExtractor (Bertin & Arnouts 1996) on the “flattened” stack, i.e., with the low spatial frequencies removed (sky background and any Galactic cirrus or nebulae). Multi-wavelength catalogues including VIS, NISP, and ground-based CFHT-MegaCam photometry have been provided in Cuillandre et al. (2024b). After star-galaxy separation and limiting the analysis to sources with , we end up with a parent sample of objects, which will be used in the visual classification procedure. As shown in Fig. 1, the adopted magnitude cut is well motivated by the fact that sources at fainter magnitudes are compact and featureless.
For the experiment, cutouts in the four bands of (i.e. pixels and pixels in the VIS and NISP bands, respectively) have been created using the “flattened” stacks.
3 Method
3.1 The visualisation tools
We use the visualisation tools prepared by Acevedo Barroso et al. (in prep.) to carry out the visual inspection.111The version used for this work is available at https://github.com/ClarkGuilty/Qt-stamp-visualizer/tree/ERO_edition The tools are based on the visualisation tools used in Savary et al. (2022) and Rojas et al. (2022), but reimplemented in the Qt 6 framework with extended functionality targeting the requirements of the Euclid ERO data. The two applications correspond to a mosaic viewer and a 1-by-1 sequential viewer, both presented in Fig. 2. The mosaic viewer displays rectangular mosaics of objects for the user to inspect. Then, the user is tasked with clicking on objects that show any signs of lensing. Additionally, the user is allowed to mark objects as “interesting, but not a lens”. This is exemplified in the left panel of Fig. 2. By contrast, the 1-by-1 sequential tool displays one object at the time and the user is tasked with classifying it into one of the following non-overlapping categories:
A: sure lens, it shows clear lensing features and no additional information is needed;
B: probable lens, it shows lensing features but additional information is required to verify that it is a definite lens;
C: possible lens, it shows lensing features, but they can be explained without resorting to gravitational lensing;
X: definitively not a lens;
Interesting: definitively not a lens, but interesting in some other way.
The grades A, B, and C are deemed as positive grades, whereas X and Interesting are deemed as negative. The 1-by-1 sequential tool is shown in the right panel of Fig. 2. Both tools allow the users to change the colour map used for the monochromatic images, as well as the function used to scale the pixel intensities. For every object to inspect, the tools generate three high-contrast images:
a monochromatic image at the VIS resolution of ;
an RGB composite image using the , , and bands, at the NISP resolution of ;
an RGB composite image using the , , and bands, at the NISP resolution of .
Before creating the composite images, we reproject the band data from the sky-coordinate system and resolution of the VIS instrument, to the corresponding ones in the NISP instrument. This aligns the images and corrects for the different pixel scales between instruments. The experts were required to use the monochromatic high-resolution band, and the composite. The composite is not shown by default in the tools, but it is also available. Additionally, the users also have access to the individual NISP bands when using the 1-by-1 sequential tool. This is shown in Fig. 2.
3.2 The visual inspections
To make better use of the large number of strong lensing experts available for the visual inspection, we split the 41 experts into two non-overlapping teams. This allows us to try different visual inspection schemes while still retaining enough experts per team for the classifications not to be dominated by noise. Team 1 comprises 22 experts, while Team 2 comprises 19. Both teams inspect the cutouts described in Sect. 2, along with six extra cutouts of simulated lenses produced following the prescription of Euclid Consortium, Metcalf et al. (in prep). The introduction of simulated mocks tests the ability of the experts to identify lenses, and serves as a sanity check of their performance. The experts are not informed about the mocks in order to avoid any biases. The six simulated lenses can be seen in Fig. 3. Both teams are given two weeks per stage. There is no communication either between or within the teams and their results are not joined until after the visual inspections are completed.
Both teams follow a two-stage approach. First, they focus on cleaning the parent sample from obvious non-lens contaminants. Then, they inspect in detail the remaining sources and settle on a final classification following the grades introduced in Sect. 3.1.
Team 1 first inspects the entire parent sample using the mosaic tool. This corresponds to cutouts after adding the mock lenses. The experts are allowed to change the number of sources per page shown in the mosaic, with most experts observing between 20 and 42 sources per mosaic page. Afterwards, Team 1 reinspects all the objects selected in the first stage using the 1-by-1 sequential tool and assigning detailed classifications.
Simultaneously, Team 2 is further divided into three groups: two groups comprising six experts each and one group comprising seven. Each group inspects one third of the sources using the 1-by-1 sequential tool, while focusing on rejecting obvious contaminants. During the first stage, no detailed classification is required. Subsequently, all the selected sources are reinspected by Team 2 as a whole, using again the 1-by-1 sequential tool but assigning detailed classifications.
Afterwards, we aggregate the final classifications of each team independently. The output is a single final classification per team for every source. Thus, if more than half of the experts voted negatively, then the source is taken as a non-lens; otherwise, we take the mode between A, B and C. In the case of a tie between A and B or A and C, A takes precedence. This prioritisation reflects the expectation that experts will only vote for A if they are confident about the presence of lensing features. If there is a tie between B and C, we favour C to minimise noise in the final B-sample. In the event of a tie among all three positive classes, B is preferred.
4 Visual inspection
The experts are anonymised at every stage of the analysis to avoid biasing the visual inspection results. We refer to the experts only by their expert ID, for example, “expert number 12”.
4.1 Team 1
The first stage of the visual inspection was completed by 20 out of the 22 experts that registered for Team 1. Figure 4 presents the number of lens candidates selected by each expert. We note that expert number 1 selected more sources for reinspection when compared to the mean and standard deviation of the other experts from Team 1, which selected sources for reinspection. Figure 5 presents the distribution of votes with and without expert number 1. We observe that the number of sources selected only by a single expert drops by half when excluding the outlier expert. Consequently, we removed the classifications from expert number 1. This allowed us to have a generous selection for the second stage, in which we included every source chosen by any of the remaining experts, while also keeping the sample reasonably small. The total number of sources selected for inspection in stage 2 was .
The second stage was completed by 17 out of the 22 experts. Figure 12 presents the number of votes per grade per expert. Similarly, Table 1 presents the average number of votes per grade for every stage that used the 1-by-1 sequential tool. Both Figs. 12 and 1 show that the experts of Team 1 were conservative when assigning the highest grade, A. This trend, however, was not the case for grades B and C, for which multiple experts assigned hundreds of votes in the sample. This indicates confusion among some experts for what constitutes a lens when the lensing features are not strikingly obvious. However, the experts with a large number of positive classifications were still a minority. Ultimately, the behaviour of all experts during the second stage was deemed acceptable and none of classifications were removed.
4.2 Team 2
First stage | Second stage | ||
Grade | Team 2 | Team 1 | Team 2 |
A | |||
B | |||
C | |||
X | |||
Total |
The first round of visual inspection was completed by 16 out of 19 experts involved in Team 2. The distribution among groups was as follows: five experts from both the first and second group, and another six from the third group. Figure 13 and Table 1 summarise the behaviour of the experts during the inspection. The grades A and B were given to only a small number of sources and the overall proportion of positive votes was low. This is consistent with the task of removing the contaminants, which was the goal for stage 1. The visual inspection results are presented in Fig. 6. We discarded all the sources with a negative majority vote, except those that got at least one vote for A or B. This amounted to sources for reinspection in stage 2.
The second stage of Team 2 was completed by 17 experts out of the 19 who registered. The grading behaviour of the experts is presented in Fig. 14. A few of the experts gave a large number of positive grades, one grading positively 673 out of the 691. Still, most of the experts reserved the grades A and B for the few best candidates, thus, we did not remove any experts. This is justified by the aggregation method introduced in Sect. 3.2, which relies on a majority vote to assign final grades.
4.3 Time cost of the inspections
It took the experts of Team 1 between and to inspect the whole parent sample using the mosaic tool, with the median time being . By contrast, it took the experts from Team 2 between to with a median time of to inspect one third of the parent sample, but using the 1-by-1 sequential tool. The median time per source was and for Teams 1 and 2, respectively.
For the second stage, Team 1 experts took between and . The median time was , equivalent to per source. Finally, Team 2 experts took between and , and the median was , which corresponds to per source. Overall, the time taken per source during the second stage was consistent for both teams. Given the filtering of the sample during the first stage, the time taken per source in the second stage corresponds more closely to the real time that a visual inspection would take on a preselected sample, e.g., the output of a CNN.
5 Results
5.1 The visual inspection candidates
We compute three grades for each source: two individual grades from Team 1 and Team 2, plus a final joint grade. The final grade of each team is determined by applying the scheme presented in Sect. 3.2 to their respective set of votes. The joint final grade is determined the same way but using the votes of both teams together in the computation, with the exception of sources that were rejected by one of the teams during the first stage. In that case, we add two negative votes to the classifications of the other team and calculate the final joint classification using the 19 votes.
Overall, we obtain three grade A, 13 grade B, and 52 grade C lens candidates from the final joint grades. We present a summary of the number of candidates per team in Fig. 7. Furthermore, we show the grade A and B candidates in Fig. 8, along with their three final grades: the final grade from Team 1, the final grade from Team 2, and the final joint grade (in yellow, cyan, and red, respectively). The grade C candidates, along with the sources selected by either team but rejected from the joint sample, are presented in Appendix B in Figs. 15, 16, 17, 18 and 19. We note that Team 1 found almost double the amount of grade A and grade B candidates, but about the same number of grade C candidates as Team 2. Moreover, all the candidates selected by Team 1 but rejected from the final sample were also rejected by Team 2 during the first stage. By contrast, most of the sources selected by Team 2 but rejected from the final sample were also rejected by Team 1 during the second stage. This suggests that using the 1-by-1 sequential tool to filter out the obvious non-lenses, as Team 2 did during the first stage, is in fact a more aggressive filter than using the mosaic tool to preselect candidates.
5.2 Modelling of the lens candidates
We further assess the validity of the 16 grade A and B candidates by modelling the lens galaxy light and mass distribution, and surface brightness distribution of the lensed source. We do this using the pronto software (Vegetti & Koopmans 2009; Rybak et al. 2015a, b; Rizzo et al. 2018; Ritondale et al. 2019; Powell et al. 2021). We use a singular isothermal ellipsoid (SIE) density profile to describe the mass distribution of the galaxy, a composite of three Sérsic profiles for the lens galaxy light, and a pixellated, free-form reconstruction of the source surface brightness distribution. The parameters for the lens light and mass models are found through a non-linear optimisation using MultiNest (Feroz et al. 2009). For each set of light and mass parameters, the surface brightness distribution of the source is solved for linearly, up to some regularisation condition, which, in this case, penalises large gradients in the source. The strength of this regularisation is itself a non-linear parameter. The full model has 23 free parameters.
We use a circular mask centred on the lensing galaxy and large enough to enclose what are assumed to be lensed images. We also use the positions of these assumed lensed images as an input to the optimisation scheme. After selecting 2, 3, or 4 positions in the image plane, the non-linear optimiser will only accept models where these image plane positions have an RMS separation in the source plane below some tolerance, in this case 1 arcsec. This removes the need to fine tune the initial modelling conditions, since only combinations of lens parameters which focus the source are accepted. The large tolerance of 1 arcsec also prevents the subjective choice of image positions from placing strict restrictions on the model.
After the optimisation, we check each model against three criteria to determine if its data are well described by a strong lensing model. These criteria are: (1) is there an SIE critical curve that can enclose or exclude the right number of bright components in the image plane; (2) is this critical curve, itself an isodensity contour, consistent with the lens galaxy light profile; and (3) is the reconstructed source surface brightness distribution consistent with a compact, focused object, inside a caustic.
Five of the 16 grade A and B candidates are found to have valid lens models according to the criteria above. Six candidates cannot be determined as lenses with the available data, and a final five candidates do not have valid lens models. For the five valid candidates, we measure Einstein radii in the range –. The result for one valid candidate is shown, as an example, in Fig. 9, with the rest being shown in Appendix C.
Name | RA | Dec | Final joint grade | Modelling | Einstein Radius [arcsec] | ||
J031715+414057 | 49.315974 | 41.682676 | 20.810 | 0.004 | B | Valid | 0.70 |
J031741+412702 | 49.424405 | 41.450783 | 22.393 | 0.007 | B | Valid | 1.01 |
J031749+420011 | 49.456529 | 42.003291 | 19.791 | 0.003 | A | Valid | 0.92 |
J031953+415746 | 49.972097 | 41.962896 | 19.576 | 0.003 | A | Valid | 1.24 |
J031959+414229 | 49.999999 | 41.708073 | 22.467 | 0.007 | B | Valid | 0.68 |
J031615+414402 | 49.063959 | 41.733984 | 21.651 | 0.005 | B | Undetermined | |
J031640+414947 | 49.169013 | 41.829837 | 21.902 | 0.007 | B | Undetermined | |
J031647+414210 | 49.196262 | 41.702880 | 21.193 | 0.003 | A | Undetermined | |
J031648+415311 | 49.200924 | 41.886438 | 21.171 | 0.005 | B | Undetermined | |
J031834+413656 | 49.642982 | 41.615827 | 21.280 | 0.002 | B | Undetermined | |
J031838+413940 | 49.659347 | 41.661254 | 20.905 | 0.002 | B | Undetermined | |
J031646+414840 | 49.195286 | 41.811343 | 20.837 | 0.002 | B | Not valid | |
J031733+414944 | 49.388333 | 41.829079 | 20.666 | 0.002 | B | Not valid | |
J031925+420547 | 49.854480 | 42.096580 | 19.461 | 0.001 | B | Not valid | |
J031930+413711 | 49.877131 | 41.619889 | 22.147 | 0.004 | B | Not valid | |
J031941+414953 | 49.924888 | 41.831457 | 22.014 | 0.003 | B | Not valid |
6 Discussion
6.1 The lens candidates in the literature
We crossmatch the sky coordinates of the 68 lens candidates against the Strong Lens Database (SLED, Vernardos et al., in prep.), a database of gravitational lenses, lens candidates, and known contaminants, encompassing more than entries. We found no matches between our sample and the database. Furthermore, the database indicates that there have been no lens searches in the area, since the closest match is more than away. This is expected given that the Perseus cluster is near the Galactic plane, and thus, is not generally targeted by large-scale optical galaxy surveys.
Nonetheless, Li et al. (2024) have performed a lens search in some of the ERO fields, including the Perseus cluster, in order to test the lens detection algorithm to be used by the Chinese Space Station Telescope (CSST) team. For that purpose, they used the high-resolution media images published by ESA.222Available here: https://www.esa.int/Science_Exploration/Space_Science/Euclid/Euclid_s_first_images_the_dazzling_edge_of_darkness They found four lens candidates in the Perseus cluster and reported the pixel coordinates in the TIFF image. Three of those candidates are included in our parent sample and were selected by Team 1 for reinspection in the second stage. However, all of them were rejected in the second stage. A smaller committee of experts inspected the fourth candidate and ultimately rejected it. Thus, we believe that none of the Li et al. (2024) candidates in the Perseus cluster are lenses. The lack of real lenses in their sample is likely a consequence of using the lower resolution TIFF images and the fact that their model was not trained to detect lenses in Euclid data. Therefore, the false positives in the Perseus cluster do not necessarily reflect on the potential performance of the algorithm in CSST data.
6.2 The expected prevalence of ELSE lenses
Collett (2015, C15 hereafter) forecasts that of Euclid imaging should contain strong gravitational lenses. Naively scaling down to the Perseus field gives an expectation of eight galaxy-galaxy lenses, of which six should have ¡ 23. We have three grade A and 13 grade B candidates, of which five have a valid lens model, and six are indeterminate. Despite the small number statistics, it is clear that our results and the C15 forecasts are broadly consistent. As well as the Poisson noise, our limited understanding of the discovery selection function, and the lack of spectroscopic redshift confirmation of our candidates makes it impossible to precisely compare our absolute number of lenses with the forecast population.
The median Einstein radius forecast in C15 is , whereas the smallest Einstein radius of the five lenses with valid models in Sect. 5.2 is . This tension is alleviated somewhat by accounting for the fact we only look at lenses with ¡ 23: by excluding the faintest lenses we prefer higher mass or lower redshift lenses, both of which result in larger Einstein radii. After applying this cut, the forecast median Einstein radius increases to . Figure 10 shows the expected number of lenses in of Euclid data, their Einstein radius distribution, and the Einstein radius distribution of successfully modelled ELSE candidates. The C15 forecasts assumed that the lens galaxy light would be subtracted, which is most helpful for finding small Einstein radius lenses: we have not done this. In future work, we will calibrate this selection function using simulations (Rojas et al., in prep.) but for now, we use a toy model to estimate it. We assume that the shape of the underlying Einstein radius population follows the Euclid forecasts of C15, but the visual inspection selection function is an additional step function where we discover every lens above a threshold Einstein radius (), and none below this threshold. Using Bayes’ Theorem, we infer given that we discovered five lenses and the lowest Einstein radius was in the range –.
We use a Monte Carlo simulation to draw many realisations of five Einstein radii, but vary the discovery threshold, assuming a uniform prior on the discovery threshold, . We find that the threshold is arcsec at 68% confidence. This corresponds to a forecast total population of lenses in the full Euclid data set that are discoverable with visual inspection without lens light subtraction and with lens ¡ 23. Moreover, scaling down the area to the forecast predicts approximately five lenses, coinciding with the five lens candidates with good models.
6.3 The prevalence of ELSE lenses
We extrapolate the five lens candidates with valid lens model found in the ERO Perseus field to the entire EWS. For this, we make three key assumptions:
1.
we have found all the discoverable lenses with within the ERO Perseus field;
2.
the average number density of sources with in the EWS is the same as in the ERO Perseus field: extended sources per ;333About of sources are cluster members and they occlude about of the sky. In terms of number of sources, the effect of the cluster balances out.
3.
the likelihood that a given source within the EWS exhibits lensing features mirrors that of the sources within the parent sample.
We define a lens as discoverable if it can be identified via expert visual inspection and a valid lens model can be found for it. This assumption means that, in fact, we are mostly estimating the prevalence of larger Einstein radius lenses, which are easier to model and experts will identify more easily. However, small Einstein radius lenses can still be identified if the source is bright and the geometry of the lensed images is particularly evident; for example, the mock lens in the bottom right panel of Fig. 3, which has an Einstein radius of , but also a typical quad geometry, was still graded A by the experts.
Under the key assumptions above, the prevalence of ELSE lenses in the EWS can be estimated as follows. The likelihood of finding lenses among trialled sources is given by the hypergeometric distribution
where and are the total numbers of lenses and sources in the EWS. Using the second assumption, we estimate the total number of extended sources with in the of the EWS to be . The number of trialled sources was , and the number of discoverable lenses was . We can then compute the posterior probability by Bayes’ theorem, i.e, . We assume that the prior probability of , , is such that the ratio is distributed uniformly in log-space between and unity. Overall, we expect ELSE-type lenses in the entire EWS. The posterior distribution is presented in Fig. 11, with the 68% confidence interval highlighted in light blue. Figure 11 also shows in orange the forecast of C15 as described in the previous section, which lies inside the confidence range of our estimate, showing a good agreement between the forecast and our estimate.
7 Conclusions
In this work we have investigated the performance of the Euclid telescope for detecting galaxy-scale gravitational lenses. We have employed 41 experts to carry out a blind visual search for lenses in the Euclid ERO data of the Perseus cluster. The search has yielded three grade A and 13 grade B lens candidates.
We have modelled the 16 candidates, to test if the observed VIS images are consistent with a single background galaxy lensed by a simple, plausible lens. We have obtained convincing models for five of the candidates. Modelling for six of the candidates produced inconclusive results, whilst five of the visual inspection candidates are definitively non-lenses based on our modelling.
We have extrapolated the five candidates with valid lens models to the entire EWS and estimated the number of lenses in it that are discoverable with visual inspection and can be confirmed through modelling of the VIS data. This extrapolation yields lenses in the full EWS. This is broadly consistent with the forecast by C15.
Even though our magnitude cut of effectively removes many of the small Einstein radii objects, the distribution of Einstein radii in our modelled sample is still on the higher side when compared to C15, indicating that we are either unable to identify small Einstein radii systems in the visual inspections, or are unable to model them under our simplifying assumptions. Assuming a step function cutoff in Einstein radius, we inferred the cutoff point to be ; below this, our methodology is likely missing most of the lens candidates. Convolving our inferred cutoff with the Collett (2015) Euclid population, we update the forecast to predict galaxy-scale lenses in the whole EWS (assuming that the same visual inspection discovery and modelling methodology are performed on the entire data set). Downscaling to the visual inspected in this work gives approximately lenses, which is in perfect accordance with the five lens candidates with valid models.
This sample represents the first gravitational lenses reported in this patch of the sky, and some of the first lenses discovered with Euclid data. There is tentative evidence that we are missing small Einstein radius lenses, so future work will be necessary to develop identification techniques targeting low-Einstein radii lens candidates. There are also substantial challenges in how we scale up from this field to the full data set: blind visual inspection of is implausible. Neural networks and citizen science are both likely to be needed to reduce the sample to a manageable size.
Even without spectroscopic confirmation of our candidates, our simple lens modelling provides compelling evidence that at least five of them are indeed strong gravitational lenses. These results are hugely promising for the future of strong lensing science with Euclid. Forecasts and early data both now point to the discovery of or more gravitational lenses in the full Euclid data set.
Acknowledgements.
This work has made use of the Early Release Observations (ERO) data from the Euclid mission of the European Space Agency (ESA), 2024, https://doi.org/10.57780/esa-qmocze3. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org). J. A. A. B., B. C., and F. C. acknowledge support from the Swiss National Science Foundation (SNSF). C. OR thanks the Max Planck Society for support through a Max Planck Lise Meitner Group. C. T. and V. B. acknowledge the INAF grant 2022 LEMON. T. E. C. is funded by a Royal Society University Research Fellowship. This project has received funding from the European Research Council (ERC) under the European Unions Horizon 2020 research and innovation programme (LensEra: grant agreement No 945536).References
- Auger et al. (2009) Auger, M. W., Treu, T., Bolton, A. S., et al. 2009, ApJ, 705, 1099
- Bertin & Arnouts (1996) Bertin, E. & Arnouts, S. 1996, A&AS, 117, 393
- Birrer et al. (2018) Birrer, S., Refregier, A., & Amara, A. 2018, ApJ, 852, L14
- Birrer et al. (2017) Birrer, S., Welschen, C., Amara, A., & Refregier, A. 2017, J. Cosmology Astropart. Phys., 2017, 049
- Canameras et al. (2023) Canameras, R., Schuldt, S., Shu, Y., et al. 2023, arXiv:2306.03136
- Collett (2015) Collett, T. E. 2015, ApJ, 811, 20
- Collett & Bacon (2016) Collett, T. E. & Bacon, D. J. 2016, MNRAS, 456, 2210
- Cuillandre et al. (2024a) Cuillandre, J.-C., Bertin, E., Bolzonella, M., et al. 2024a, A&A, submitted, arXiv:2405.13496
- Cuillandre et al. (2024b) Cuillandre, J.-C., Bolzonella, M., Boselli, A., et al. 2024b, A&A, submitted, arXiv:2405.13501
- Dutton & Treu (2014) Dutton, A. A. & Treu, T. 2014, MNRAS, 438, 3594
- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, submitted, arXiv:2405.13492
- Euclid Collaboration: Jahnke et al. (2024) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2024, A&A, submitted, arXiv:2405.13493
- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., Achúcarro, A., et al. 2024, A&A, submitted, arXiv:2405.13491
- Euclid Collaboration: Scaramella et al. (2022) Euclid Collaboration: Scaramella, R., Amiaux, J., Mellier, Y., et al. 2022, A&A, 662, A112
- Euclid Early Release Observations (2024) Euclid Early Release Observations. 2024, https://doi.org/10.57780/esa-qmocze3
- Faure et al. (2008) Faure, C., Kneib, J.-P., Covone, G., et al. 2008, ApJS, 176, 19
- Feroz et al. (2009) Feroz, F., Hobson, M. P., & Bridges, M. 2009, MNRAS, 398, 1601
- Ferreras et al. (2010) Ferreras, I., Saha, P., Leier, D., Courbin, F., & Falco, E. E. 2010, MNRAS, 409, L30
- Fleury et al. (2021) Fleury, P., Larena, J., & Uzan, J.-P. 2021, J. Cosmology Astropart. Phys., 2021, 024
- Garvin et al. (2022) Garvin, E. O., Kruk, S., Cornen, C., et al. 2022, A&A, 667, A141
- Gilman et al. (2024) Gilman, D., Birrer, S., Nierenberg, A., & Oh, M. S. H. 2024, MNRAS[arXiv:2403.03253]
- Grespan et al. (2024) Grespan, M., Thuruthipilly, H., Pollo, A., et al. 2024, arXiv:2405.11992
- Grillo et al. (2024) Grillo, C., Pagano, L., Rosati, P., & Suyu, S. H. 2024, A&A, 684, L23
- Hezaveh et al. (2013) Hezaveh, Y. D., Marrone, D. P., Fassnacht, C. D., et al. 2013, ApJ, 767, 132
- Hogg et al. (2023) Hogg, N. B., Fleury, P., Larena, J., & Martinelli, M. 2023, MNRAS, 520, 5982
- Jacobs et al. (2019) Jacobs, C., Collett, T., Glazebrook, K., et al. 2019, ApJS, 243, 17
- Keeton et al. (2000) Keeton, C. R., Mao, S., & Witt, H. J. 2000, ApJ, 537, 697
- Li et al. (2024) Li, X., Sun, R., Lv, J., et al. 2024, AJ, 167, 264
- Marshall et al. (2015) Marshall, P. J., Lintott, C. J., & Fletcher, L. N. 2015, ARA&A, 53, 247
- More et al. (2016) More, A., Verma, A., Marshall, P. J., et al. 2016, MNRAS, 455, 1191
- Oguri & Marshall (2010) Oguri, M. & Marshall, P. J. 2010, MNRAS, 405, 2579
- Orban de Xivry & Marshall (2009) Orban de Xivry, G. & Marshall, P. 2009, MNRAS, 399, 2
- O’Riordan et al. (2023) O’Riordan, C. M., Despali, G., Vegetti, S., Lovell, M. R., & Moliné, Á. 2023, MNRAS, 521, 2342
- Paraficz et al. (2018) Paraficz, D., Rybak, M., McKean, J. P., et al. 2018, A&A, 613, A34
- Pascale et al. (2024) Pascale, M., Frye, B. L., Pierel, J. D. R., et al. 2024, arXiv:2403.18902
- Pawase et al. (2014) Pawase, R. S., Courbin, F., Faure, C., Kokotanekova, R., & Meylan, G. 2014, MNRAS, 439, 3392
- Petrillo et al. (2019) Petrillo, C. E., Tortora, C., Chatterjee, S., et al. 2019, MNRAS, 482, 807
- Powell et al. (2021) Powell, D., Vegetti, S., McKean, J. P., et al. 2021, MNRAS, 501, 515
- Refsdal (1964) Refsdal, S. 1964, MNRAS, 128, 307
- Ritondale et al. (2019) Ritondale, E., Vegetti, S., Despali, G., et al. 2019, MNRAS, 485, 2179
- Rizzo et al. (2018) Rizzo, F., Vegetti, S., Fraternali, F., & Di Teodoro, E. 2018, MNRAS, 481, 5606
- Rojas et al. (2023) Rojas, K., Collett, T. E., Ballard, D., et al. 2023, MNRAS, 523, 4413
- Rojas et al. (2022) Rojas, K., Savary, E., Clément, B., et al. 2022, A&A, 668, A73
- Rybak et al. (2015a) Rybak, M., McKean, J. P., Vegetti, S., Andreani, P., & White, S. D. M. 2015a, MNRAS, 451, L40
- Rybak et al. (2015b) Rybak, M., Vegetti, S., McKean, J. P., Andreani, P., & White, S. D. M. 2015b, MNRAS, 453, L26
- Savary et al. (2022) Savary, E., Rojas, K., Maus, M., et al. 2022, A&A, 666, A1
- Shajib et al. (2023) Shajib, A. J., Mozumdar, P., Chen, G. C. F., et al. 2023, A&A, 673, A9
- Sonnenfeld et al. (2019) Sonnenfeld, A., Jaelani, A. T., Chan, J., et al. 2019, A&A, 630, A71
- Stein et al. (2022) Stein, G., Blaum, J., Harrington, P., Medan, T., & Lukić, Z. 2022, ApJ, 932, 107
- Thuruthipilly et al. (2022) Thuruthipilly, H., Zadrozny, A., Pollo, A., & Biesiada, M. 2022, A&A, 664, A4
- Vegetti & Koopmans (2009) Vegetti, S. & Koopmans, L. V. E. 2009, MNRAS, 392, 945
- Vegetti et al. (2010) Vegetti, S., Koopmans, L. V. E., Bolton, A., Treu, T., & Gavazzi, R. 2010, MNRAS, 408, 1969
- Wong et al. (2020) Wong, K. C., Suyu, S. H., Chen, G. C. F., et al. 2020, MNRAS, 498, 1420
Appendix A Expert classifications in visual inspections
We show in this appendix the detailed classifications given by the experts when using the 1-by-1 sequential tool.
Appendix B Other lens candidates
In this appendix, we present the lens candidates discovered during the visual search, but that did not become part of our main sample. These samples are dominated by contaminants, making them more representative of the typical false positives found in lens searches rather than of good lens candidates.
Appendix C Modelling results
The results of modelling all 16 grade A and B candidates are shown in Fig. 20.